### v2

@@ Abstract
Keywords:
Large Language Models Knowledge Graph Multi-Agent
@@ 1 Introduction
Knowledge graphs (KGs) [1] represent facts in the real world as triples, such as (Paris, capital_of, France), facilitating the organization and scaling of information [5], and have gained paramount importance in knowledge-based systems, such as retrieval-augmented generation and recommendation systems [20]. However, most large-scale KGs [1], built using rule-based methods and statistics-based methods to extract web data, often contain noisy or incorrect triples. For instance, the widely used knowledge graph NELL [15] contains around 600K incorrect triples, which account for 26% of the set of triples in NELL. Most knowledge graph-driven tasks assume all triples are correct [15], overlooking the impact of errors, which significantly degrades the performance of downstream tasks. This highlights the urgent need for effective KG error detection methods.
Existing methods can be broadly categorized into two types based on their evidence utilization. First, explicit evidence-based methods extract paths or subgraphs from the knowledge graph that support the truth of triples and use these explicit evidences to predict the credibility of triples [16]. Second, embedding-based methods predict the credibility of triples by embedding entities and relations into vector space and calculating the embedded representation of paths or subgraphs [9, 7]. Moreover, when combined with contrastive learning and pre-trained models, their performance has shown significant improvement [24, 13, 14].
@@ 2 Related Work
Knowledge Graph Error Detection: Knowledge graph error detection [15] includes methods based on explicit and implicit evidence.
Explicit Evidence : These methods evaluate the credibility of triples using direct evidence from the graph, such as paths and subgraphs. For example: Shi et al. [16] used graph traversal to extract meta-paths and calculate path weights. Lin et al. [12] applied graph pattern mining with metrics like support and confidence. However, these methods may struggle with incomplete knowledge graphs and underutilize contextual information.
Implicit Evidence : These methods use embedding models to map entities and relations into vector spaces for credibility evaluation. CKRL [21] optimized triple credibility by considering local and global factors during embedding learning. KGTtm [9] combined random walks and embeddings, using MLP for scoring. Recent methods also use pre-trained language models and contrastive learning: KG-BERT [24] enriches semantic representations with pre-trained models. SeSICL [14], StAR [19], and CCA [13] align textual and structural representations via contrastive learning. These methods improve performance but rely on fixed graph structures, limiting adaptability to new or unseen structures.
Multiple LLM-based Agent System: Multi-Agent systems such as CAMEL  [10], have gained significant attention for simulating complex interactions among intelligent agents. These systems coordinate multiple LLM agents to tackle complex tasks, particularly in NLP evaluation. For example, ChatEval [3] uses a multi-agent jury to explore the impact of communication strategies on evaluating open-ended questions and traditional NLG tasks, while MATEval [11] involves agents collaborating to evaluate story texts, enhancing the evaluation process through multi-agent interaction.
@@ 3.1 Knowledge Graph Error
A triple is incorrect if the head or tail entity does not align with the relation [22]. For example, (Harvard University, is_located_in, New York). Conversely, a triple is correct if all components align appropriately.
@@ 5.1 Experimental Settings
Datasets: We use two real-world knowledge graph datasets: FB15K [17] and WN18RR [6]. We chose these two datasets because they are highly representative in the field of knowledge graph error detection, encompassing most typical scenarios and possible graph structural representations found in knowledge graph data. In each dataset, we simulate realistic errors by replacing entities and relations with similar ones selected based on cosine similarity within the dataset, resulting in approximately 30% of the data being erroneous. We split each dataset into training, validation, and test sets with a ratio of 8:1:1. The fine-tuning process used only the training set, while the test set was reserved for final evaluations. FB15K is derived from Freebase and contains a rich set of entities and relations, while WN18RR is a subset of WordNet with corrected inverse relations, increasing the complexity. Additionally, we conducted experiments on a knowledge graph dataset from China Mobile’s business scenarios, achieving the best results compared to other methods.
Baselines: We compare MAKGED against various baseline methods, including traditional knowledge graph embedding models such as TransE [2], DistMult [23], and ComplEx [18], which learn triple embeddings to compute confidence scores. Additionally, we compared recent embedding-based KG error detection methods including CAGED [25] and KGTtm [9]. We also compared KG-BERT [24] and CSProm-KG [4] models that combine pre-trained language models for error detection, as well as models that use text structure and graph structure for comparative learning, such as StAR [19], SeSICL [14], and CCA [13] models. All comparison experiments are conducted under the same experimental settings.
Implementation Details: We use the Llama2 for fine-tuning and employ LoRA [8] for instruction tuning. Experiments are conducted on V100 GPU servers. GCN Component: A three-layer GCN with 128 hidden dimensions and 64-dimensional embeddings is trained separately using the Adam optimizer (learning rate: 0.001), a batch size of 64, for 100 epochs. The resulting GCN embeddings are concatenated with Llama2’s text embeddings to create a unified representation for each agent. Fine-Tuning Llama2: During fine-tuning, the combined embeddings (GCN + Llama2) serve as model inputs. We utilize mixed precision training and gradient checkpointing to accelerate training and reduce memory usage. We use Accuracy, F1-Score, Precision, and Recall as evaluation metrics. These metrics use macro averaging for both classes. Each round of agent discussion took an average of 2.3 seconds. This demonstrates the framework’s practical feasibility for industrial applications.

### v5

@@ 1 Introduction
Knowledge graphs (KGs) [kg] represent facts in the real world as triples, such as (Paris, capital_of, France), facilitating the organization and scaling of information [kg-info], and have gained paramount importance in knowledge-based systems, such as retrieval-augmented generation and recommendation systems [recommend]. However, most large-scale KGs [kg], built using rule-based methods and statistics-based methods to extract web data, often contain noisy or incorrect triples. For instance, the widely used knowledge graph NELL [survey] contains around 600K incorrect triples, which account for 26% of the set of triples in NELL. Most knowledge graph-driven tasks assume all triples are correct [survey], overlooking the impact of errors, which significantly degrades the performance of downstream tasks. This highlights the urgent need for effective KG error detection methods.
Existing methods can be broadly categorized into two types based on their evidence utilization. First, explicit evidence-based methods extract paths or subgraphs from the knowledge graph that support the truth of triples and use these explicit evidences to predict the credibility of triples [shi]. Second, embedding-based methods predict the credibility of triples by embedding entities and relations into vector space and calculating the embedded representation of paths or subgraphs [kgttm, embedding2]. Moreover, when combined with contrastive learning and pre-trained models, their performance has shown significant improvement [kg-bert, cca, sesicl].
@@ 2 Related Work
Knowledge Graph Error Detection: Knowledge graph error detection [survey] includes methods based on explicit and implicit evidence.
Explicit Evidence : These methods evaluate the credibility of triples using direct evidence from the graph, such as paths and subgraphs. For example: Shi et al. [shi] used graph traversal to extract meta-paths and calculate path weights. Lin et al. [lin] applied graph pattern mining with metrics like support and confidence. However, these methods may struggle with incomplete knowledge graphs and underutilize contextual information.
Implicit Evidence : These methods use embedding models to map entities and relations into vector spaces for credibility evaluation. CKRL [CKRL] optimized triple credibility by considering local and global factors during embedding learning. KGTtm [kgttm] combined random walks and embeddings, using MLP for scoring. Recent methods also use pre-trained language models and contrastive learning: KG-BERT [kg-bert] enriches semantic representations with pre-trained models. SeSICL [sesicl], StAR [star], and CCA [cca] align textual and structural representations via contrastive learning. These methods improve performance but rely on fixed graph structures, limiting adaptability to new or unseen structures.
Multiple LLM-based Agent System: Multi-Agent systems such as CAMEL  [camel], have gained significant attention for simulating complex interactions among intelligent agents. These systems coordinate multiple LLM agents to tackle complex tasks, particularly in NLP evaluation. For example, ChatEval [chateval] uses a multi-agent jury to explore the impact of communication strategies on evaluating open-ended questions and traditional NLG tasks, while MATEval [mateval] involves agents collaborating to evaluate story texts, enhancing the evaluation process through multi-agent interaction.
@@ 3.1 Knowledge Graph Error
A triple is incorrect if the head or tail entity does not align with the relation [error]. For example, (Harvard University, is_located_in, New York). Conversely, a triple is correct if all components align appropriately.
@@ 5.1 Experimental Settings
Datasets: We use two real-world knowledge graph datasets: FB15K [fb15k] and WN18RR [wn18rr]. We chose these two datasets because they are highly representative in the field of knowledge graph error detection, encompassing most typical scenarios and possible graph structural representations found in knowledge graph data. In each dataset, we simulate realistic errors by replacing entities and relations with similar ones selected based on cosine similarity within the dataset, resulting in approximately 30% of the data being erroneous. We split each dataset into training, validation, and test sets with a ratio of 8:1:1. The fine-tuning process used only the training set, while the test set was reserved for final evaluations. FB15K is derived from Freebase and contains a rich set of entities and relations, while WN18RR is a subset of WordNet with corrected inverse relations, increasing the complexity. Additionally, we conducted experiments on a knowledge graph dataset from China Mobile’s business scenarios, achieving the best results compared to other methods.
Baselines: We compare MAKGED against various baseline methods, including traditional knowledge graph embedding models such as TransE [transe], DistMult [distmult], and ComplEx [complex], which learn triple embeddings to compute confidence scores. Additionally, we compared recent embedding-based KG error detection methods including CAGED [CAGED] and KGTtm [kgttm]. We also compared KG-BERT [kg-bert] and CSProm-KG [csprom] models that combine pre-trained language models for error detection, as well as models that use text structure and graph structure for comparative learning, such as StAR [star], SeSICL [sesicl], and CCA [cca] models. All comparison experiments are conducted under the same experimental settings.
Implementation Details: We use the Llama2 for fine-tuning and employ LoRA [lora] for instruction tuning. Experiments are conducted on V100 GPU servers. GCN Component: A three-layer GCN with 128 hidden dimensions and 64-dimensional embeddings is trained separately using the Adam optimizer (learning rate: 0.001), a batch size of 64, for 100 epochs. The resulting GCN embeddings are concatenated with Llama2’s text embeddings to create a unified representation for each agent. Fine-Tuning Llama2: During fine-tuning, the combined embeddings (GCN + Llama2) serve as model inputs. We utilize mixed precision training and gradient checkpointing to accelerate training and reduce memory usage. We use Accuracy, F1-Score, Precision, and Recall as evaluation metrics. These metrics use macro averaging for both classes. Each round of agent discussion took an average of 2.3 seconds. This demonstrates the framework’s practical feasibility for industrial applications.
@@ 6 Conclusion
7 Acknowledgments
This work is partially supported by National Nature Science Foundation of China under No. U21A20488, and is funded by Southeast University-China Mobile Research Institute Joint Innovation Center. We thank the Big Data Computing Center of Southeast University for providing the facility support on the numerical calculations in this paper.